OpenStudy (anonymous):

i have 2x^2y/3n*2xn^2=4x^3/6n^3x is this correct

OpenStudy (anonymous):

its multiplying fort his one

OpenStudy (agent0smith):

\[\huge \frac{2x^2 y }{3n }*2xn^2\] ?

OpenStudy (anonymous):

yes

OpenStudy (agent0smith):

Then your answer isn't correct, it's missing a y for a start

OpenStudy (agent0smith):

\[\huge \frac{2x^2 y*2xn^2 }{3n } \]simplify this way.

OpenStudy (anonymous):

oh, 4n^2x^3y

OpenStudy (anonymous):

4n^2x^3y/6xn^3

OpenStudy (agent0smith):

4n^2x^3y isn't correct

OpenStudy (anonymous):

oh i see i got 4n^3x^3y/3

OpenStudy (agent0smith):

Not right... what is n^2/n?

OpenStudy (anonymous):

n

OpenStudy (agent0smith):

see your error?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

4x^3y/3

OpenStudy (agent0smith):

not quite... you already said n^2/n equals n, now you're missing an n\[\huge \frac{2x^2 y*2xn^2 }{3n }\]

OpenStudy (anonymous):

4x^3yn/3

OpenStudy (agent0smith):

yes

OpenStudy (anonymous):

thanks